# Small maximal clusters are very unlikely in critical random graphs

### v1

@@ Abstract
We describe a probabilistic methodology, based on random walk estimates, to obtain exponential upper bounds for the probability of observing unusually small maximal components in two classical (near-)critical random graph models. More specifically, we analyse the near-critical Erdős-Rényi model $\mathbb{G}(n,p)$ and the random graph $\mathbb{G}(n,d,p)$ obtained by performing near-critical $p$-bond percolation on a simple random $d$-regular graph and show that, for each one of these models, the probability that the size of a largest component is smaller than $n^{2/3}/A$ is at most of order $\exp(-A^{3/2})$. The exponent $3/2$ is known to be optimal for the near-critical $\mathbb{G}(n,p)$ random graph, whereas for the near-critical $\mathbb{G}(n,d,p)$ model the best known upper bound for the above probability was of order $A^{-3/5}$. As a secondary result we show, by means of an optimized version of the martingale method of Nachmias and Peres, that the above probability of observing an unusually small maximal component is at most of order $\exp(-A^{3/5})$ in other two critical models, namely a random intersection graph and the quantum random graph; this stretched-exponential bounds also improve upon the known (polynomial) bounds available for these other two critical models.
@@ 1 Introduction
In recent years, the problem of deriving exponential bounds for the probability of observing unusually large maximal components in critical random graphs has gained considerable interest; see e.g. [30, 22, 31, 12, 13] as well as Section 7.1.1 in the PhD thesis of Dhara [16].
@@
where $\mathcal{C}_{\max}(\mathbb{G})$ is a maximal cluster of some (critical) random graph $\mathbb{G}$ and $k\in\mathbb{N}$ (for us, $k=n^{2/3}/A$, with $n^{2/3}$ being the correct order of $|\mathcal{C}_{\max}(\mathbb{G})|$ in the regime we are interested in).
@@
This random graph, which is obtained by performing $p$-bond percolation on the complete graph on $n$ vertices (i.e. it is constructed by independently retaining each edge with probability $p$ and deleting it with probability $1-p$), is known to undergo a phase transition as $np$ passes 1. Specifically, letting $p=p(n)\coloneqq\mu/n$, if $\mu<1$ then $|\mathcal{C}_{\max}(\mathbb{G}(n,p))|$ is of order $\log n$; if $\mu=1$ (the critical case), then $|\mathcal{C}_{\max}(\mathbb{G}(n,p))|$ is of order $n^{2/3}$; and if $\mu>1$, then $|\mathcal{C}_{\max}(\mathbb{G}(n,p))|$ is of order $n$. See e.g. the monographs [7], [20] or [24] for more details.
Actually we know more: if $\mu=\mu(n)=1+\lambda n^{-1/3}$ then $|\mathcal{C}_{\max}|$ is still of order $n^{2/3}$ [20]; this is the so-called critical window. Pittel [30] considered this regime and he showed (among many other things) that, in the near-critical Erdős-Rényi random graph $\mathbb{G}(n,p)$ with $p=p(n)=(1+\lambda n^{-1/3})/n$ and $\lambda\in\mathbb{R}$, as $A\rightarrow\infty$
@@
where $c_{1},c_{2}$ and $c_{3}$ are positive constants (independent of $\lambda$) and $C(\lambda)$ is a positive constant which depends on $\lambda$. We refer the reader to Corollary $2$ in Pittel’s paper [30] for the definition (and numerical estimates) of the constants appearing in (2).
To the best of our knowledge, with the exception of (2), all other upper bounds available in the literature for the probability of observing small maximal clusters in (near-)critical random graphs are of the type $A^{-\gamma}$, where $\gamma\in(0,1)$.
@@ 1.1 Main result
Pittel’s proof of (2) heavily relies on combinatorial arguments and seems difficult to be adapted to other models. On the other hand, the probabilistic argument of Nachmias and Peres [28, 27], although very general and robust, does not seem sufficient to retrieve the optimal exponent $3/2$ which appears in the leading term of (2). Indeed, although we can boost (by means of few modifications) the original martingale argument [28, 27] to obtain upper bounds of the type $\exp(-A^{3/5})$ in several critical models, it seems that the method can’t be pushed to achieve upper bounds of the (correct) order $\exp(-A^{3/2})$; we refer the reader to Section 3 for a detailed explanation concerning why we believe this to be the case.
The goal of this work is to introduce a robust probabilistic methodology, simply based on random walk estimates, to derive sharp exponential upper bounds for the probability of observing unusually small maximal clusters in (near-)critical random graphs. We do so by analysing the near-critical models $\mathbb{G}(n,p)$ and $\mathbb{G}(n,d,p)$ and show that $|\mathcal{C}_{\max}(\mathbb{G})|\geq n^{2/3}/A$ with probability at least of order $1-e^{-A^{3/2}}$ when either $\mathbb{G}=\mathbb{G}(n,p)$ or $\mathbb{G}=\mathbb{G}(n,d,p)$. Here is our main result.
@@
Let $p=(1+\lambda n^{-1/3})/n$ with $\lambda\in\mathbb{R}$ fixed. There exist constants $A_{0}=A_{0}(\lambda)>0$ and $c>0$ such that, for any $A_{0}\leq A=o(n^{2/3})$ and for all large enough $n$, we have
@@
Let $p=(1+\lambda n^{-1/3})/(d-1)$ with $\lambda\in\mathbb{R},d\geq 3$ both fixed. There exist constants $A_{0}=A_{0}(\lambda,d)>0$ and $c=c(d)>0$ such that, for any $A_{0}\leq A=O(n^{1/6})$ and for all large enough $n$, we have
@@
We emphasizes that (5) is not new (see (2) above), but our proof is purely probabilistic; on the other hand, (6) is new and most likely is of the right order; moreover, it substantially improves upon the known bound displayed in (3). We remark that Theorem 1.1 only considers the near-critical models $\mathbb{G}(n,p)$ and $\mathbb{G}(n,d,p)$, but we believe that the argument used to establish (5) and (6) is robust. One could try to adapt the methodology introduced in this work to prove similar tail bounds in other (near-)critical random graphs. We refer the reader to Section 3 for a detailed discussion (and related results) along these lines.
@@
We emphasize that, in Theorem 1.1, $A$ is allowed to depend on $n$. The condition $A=O(n^{1/6})$, which we require in our upper bound for the $\mathbb{G}(n,d,p)$ random graph could be changed to $A=o(n^{2/3})$ at the cost of adding a term of order $\exp(-A^{1/2}n^{1/6})$ to (6). Moreover, the condition $A=o(n^{2/3})$ which we require in our upper bound for the $\mathbb{G}(n,p)$ model could be relaxed to $A\leq\varepsilon n^{2/3}$ for a sufficiently small constant $\varepsilon>0$. However, when $A\asymp n^{2/3}$, then $n^{2/3}/A\asymp 1$ and so we would be bounding from above the probability that the size of a maximal cluster is at most some constant, which is not very interesting. Therefore we can say that our upper bound in (5) concerns all the relevant values of $A$. On the other hand, it could be possible to weaken the requirement $A=O(n^{1/6})$ which we made for the $\mathbb{G}(n,d,p)$ random graph at the expense of longer computations, but we refrained to do so. Indeed, we believe the bounds in (5) and (6) to be more interesting for ‘small’ (possibly dependent of $n$) values of $A$, being $n^{2/3}$ the correct order for the size of a largest cluster in both (near-)critical models. We also remark that the parameter $\lambda$, which we consider as fixed, could be allowed to depend on $n$ at the expense of more careful computations. In this case, one would need to keep track of the dependence between $A$ and $\lambda$. Furthermore, if needed, one could give bounds on the ‘smallest values’ of $A_{0}$ for which (5) and (6) hold; similarly, one might try to give bounds on the smallest $n$ for which (5) and (6) are verified (as it is done e.g. in [28]). However, we refrained to do so in order to keep the computations shorter.
@@ 1.2 Related work
They also considered the near-critical model where $p=(1+\lambda n^{-1/3})/n$ for fixed $\lambda\in\mathbb{R}$ and established a similar bound.
@@
Amongst other results (some of which were discussed earlier), Nachmias and Peres [27] proved that, in the near-critical regime where $p=(1+\lambda n^{-1/3})/(d-1)$ and $\lambda\in\mathbb{R},d\geq 3$ are both fixed, there are positive constants $c=c(\lambda,d)$ and $C=C(\lambda,d)$ which depend on both $\lambda$ and $d$ such that, for any $A>0$ and for all large enough $n$,
@@
With the purpose of introducing a probabilistic methodology capable of recovering the correct asymptotic of Pittel displayed in (7), more recently the author and Roberts [12] introduced a new methodology, based on (probabilistic) tools such as generalised ballot theorems [1] and strong embedding with Brownian motion [8], to derive matching upper and lower bounds for the probability of observing unusually large maximal clusters in critical random graphs.
In particular, for the near-critical $\mathbb{G}(n,p)$ model with $p=(1+\lambda n^{-1/3})/n$ and $\lambda\in\mathbb{R}$ (possibly dependent on $n$), the authors showed in [12] that there exist constants $A_{0}>0,n_{0}\in\mathbb{N}$ such that, if $A_{0}\leq A=A(n)=o(n^{1/30})$ and $\lambda=\lambda(n)$ satisfies $|\lambda|\leq A/3$, then for all $n\geq n_{0}$
$\mathbb{P}(|\mathcal{C}_{\max}(\mathbb{G}(n,p))|>An^{2/3})\asymp\frac{c_{2}}{A^{3/2}}\exp\big{(}-\frac{A^{3}}{8}+\frac{\lambda A^{2}}{2}-\frac{\lambda^{2}A}{2}\big{)}$ (10)
(where we use the symbol $\asymp$ to mean that the expression on the left-hand side is bounded from above and below by a constant times the expression on the right-hand side), thus improving upon the Nachmias and Peres bound given in (8) and retrieving the asymptotic result of Pittel displayed in (7).
Subsequently, the methodology which lead to (10) was adapted by the same authors to study the near-critical $\mathbb{G}(n,d,p)$ random graph where $p=(1+\lambda n^{-1/3})/(d-1)$ with $\lambda\in\mathbb{R}$ (possibly dependent on $n$) and $d\geq 3$ fixed. For this model, it was shown in [13] that there exist constants $A_{0}>0,n_{0}\in\mathbb{N}$ such that, if $A_{0}\leq A=A(n)=o(n^{1/30})$ and $\lambda=\lambda(n)$ satisfies $|\lambda|\leq A(1-2/d)[3(d-1)]^{-1}$, then for all $n\geq n_{0}$
$\mathbb{P}(|\mathcal{C}_{\max}(\mathbb{G}(n,d,p))|>An^{2/3})\asymp\frac{c_{2}}{A^{3/2}}\exp\Big{(}-\frac{A^{3}(d-1)(d-2)}{8d^{2}}+\frac{\lambda A^{2}(d-1)}{2d}-\frac{\lambda^{2}A(d-1)}{2(d-2)}\Big{)},$
@@
Observe that, in the particular case where $\lambda=0$ (that is, in the purely critical regime), the constant $c(d)\coloneqq(d-1)(d-2)/d^{2}$ which multiplies the leading term $A^{3}/8$ in the exponential above tends to $1$ as $d\rightarrow\infty$. We suspect that this phenomenon is true more generally. Specifically, let $\mathbb{G}$ be a random graph that, when considered at criticality of its parameter(s), exhibits maximal components containing $\asymp n^{2/3}$ vertices. We believe that, at criticality, there exist constants $c_{1},c_{2}>0$ with $c_{1}<c_{2}$ such that, for all large enough $n$ and (not ‘too large’) $A=A(n)$,
$\mathbb{P}(|\mathcal{C}_{\max}(\mathbb{G})|>An^{2/3})\asymp A^{-3/2}e^{-C^{*}\frac{A^{3}}{8}},$

### v3

@@ Abstract
We describe a probabilistic methodology, based on random walk estimates, to obtain exponential upper bounds for the probability of observing unusually small maximal components in two classical (near-)critical random graph models. More specifically, we analyse the near-critical Erdős-Rényi model $\mathbb{G}(n,p)$ and the random graph $\mathbb{G}(n,d,p)$ obtained by performing near-critical $p$-bond percolation on a simple random $d$-regular graph and show that, for each one of these models, the probability that a largest component contains less than $n^{2/3}/A$ vertices is at most $\exp(-\Omega(A^{3/2}))$. The exponent $3/2$ is known to be optimal for the near-critical $\mathbb{G}(n,p)$ random graph, whereas for the near-critical $\mathbb{G}(n,d,p)$ model the best known upper bound for the above probability was of (polynomial) order $O(A^{-3/5})$. As a secondary result we show, by means of an optimized version of the martingale method of Nachmias and Peres, that the above probability of observing an unusually small maximal component is at most $\exp(-\Omega(A^{3/5}))$ in other two critical models, namely a random intersection graph and the quantum random graph. These stretched-exponential bounds also improve upon the known (polynomial) upper bounds available for these other two critical models.
@@ 1 Introduction
In recent years, the problem of deriving exponential bounds for the probability of observing unusually large maximal components in critical random graphs has gained increasing interest; see e.g. [30, 22, 31, 12, 13] as well as Section 7.1.1 in the PhD thesis of Dhara [16].
@@
where $\mathcal{C}_{\max}(\mathbb{G})$ is a maximal cluster of some (near-)critical random graph $\mathbb{G}$ and $k\in\mathbb{N}$ (for us, $k=\Theta(n^{2/3}/A)$, with $n^{2/3}$ being the correct order of $|\mathcal{C}_{\max}(\mathbb{G})|$ in the regime we are interested in for the random graphs under examination).
@@
This random graph, which is obtained by performing $p$-bond percolation on the complete graph on $n$ vertices (which means that it is constructed by independently retaining each edge with probability $p$ and deleting it with probability $1-p$), is known to undergo a phase transition as $np$ passes 1. Specifically, letting $p=p(n)\coloneqq\mu/n$, the following statements hold with probability tending to one as the number of vertices grows indefinitely: if $\mu<1$ then $|\mathcal{C}_{\max}(\mathbb{G}(n,p))|$ is $O(\log n)$; if $\mu=1$ (the critical case), then $|\mathcal{C}_{\max}(\mathbb{G}(n,p))|$ is $\Theta(n^{2/3})$; and if $\mu>1$, then $|\mathcal{C}_{\max}(\mathbb{G}(n,p))|$ is $\Theta(n)$. See for instance the monographs [7], [20] or [24] for more details.
Actually we know more: if $\mu=\mu(n)=1+\lambda n^{-1/3}$ then $|\mathcal{C}_{\max}|$ is still $\Theta(n^{2/3})$ [20]; this is the so-called critical window. Pittel [30] considered this regime and he showed (among many other things) that, in the near-critical Erdős-Rényi random graph $\mathbb{G}(n,p)$ with $p=p(n)=(1+\lambda n^{-1/3})/n$ and $\lambda\in\mathbb{R}$, as $A\rightarrow\infty$
@@
where $c_{1},c_{2}$ and $c_{3}$ are positive constants (independent of $\lambda$) and $C(\lambda)$ is a positive constant which depends on $\lambda$. The asymptotic in (2) shows that, for large $A$ and $n$,
$-\log\Big{(}\mathbb{P}(|\mathcal{C}_{\max}(\mathbb{G}(n,p))|<n^{2/3}/A)\Big{)}=\Theta(A^{3/2}),$
which we also write as
$\mathbb{P}(|\mathcal{C}_{\max}(\mathbb{G}(n,p))|<n^{2/3}/A)=\exp(-\Theta(A^{3/2})).$
We refer the reader to Corollary $2$ in Pittel’s paper [30] for the definition (and numerical estimates) of the constants appearing in (2).
To the best of our knowledge, with the exception of (2), all other upper bounds available in the literature for the probability of observing small maximal clusters in (near-)critical random graphs are of much weaker order, namely the type $A^{-\gamma}$, where $\gamma\in(0,1)$.
@@ 1.1 Main result
Pittel’s proof of (2) heavily relies on combinatorial arguments and seems difficult to be adapted to other models. On the other hand, the probabilistic argument of Nachmias and Peres [28, 27], although very general and robust, does not seem sufficient to retrieve the optimal exponent $3/2$ which appears in the leading term of (2). Indeed, although we can boost (by means of a few modifications) the original martingale argument [28, 27] to obtain upper bounds of the type $\exp(-\Omega(A^{3/5}))$ in several critical models, it seems that the method can’t be pushed to achieve the (correct) decay $\exp(-\Omega(A^{3/2}))$, even for the critical $\mathbb{G}(n,p)$ model. We refer the reader to Section 3 for a detailed explanation concerning why we believe this to be the case.
The goal of this work is to introduce a robust probabilistic methodology, simply based on random walk estimates, to derive sharp exponential upper bounds for the probability of observing unusually small maximal clusters in (near-)critical random graphs. We do so by analyzing the near-critical models $\mathbb{G}(n,p)$ and $\mathbb{G}(n,d,p)$ and show that $|\mathcal{C}_{\max}(\mathbb{G})|<n^{2/3}/A$ with probability at most $\exp(-\Omega(A^{3/2}))$ when either $\mathbb{G}=\mathbb{G}(n,p)$ or $\mathbb{G}=\mathbb{G}(n,d,p)$. Here is our main result.
@@
Let $p=p(n)\coloneqq(1+\lambda n^{-1/3})/n$ with $\lambda\in\mathbb{R}$ fixed. There exist constants $A_{0}=A_{0}(\lambda)>0$ and $c>0$ such that, for any $A_{0}\leq A=A(n)=o(n^{2/3})$ and for all large enough $n$, we have
@@
Let $p=p(d,n)\coloneqq(1+\lambda n^{-1/3})/(d-1)$ with $\lambda\in\mathbb{R},3\leq d\in\mathbb{N}$ both fixed. There exist constants $A_{0}=A_{0}(\lambda,d)>0$ and $c=c(d)>0$ such that, for any $A_{0}\leq A=A(n)=O(n^{1/6})$ and for all large enough $n$ (with $dn$ even), we have
@@
We emphasizes once more that (5) is not new (see (2) above), but our proof is purely probabilistic. On the other hand, (6) is new and most likely captures the correct decay; moreover, it substantially improves upon the known bound displayed in (3). We remark that Theorem 1.1 only considers the near-critical models $\mathbb{G}(n,p)$ and $\mathbb{G}(n,d,p)$, but we believe that the argument used to establish (5) and (6) is robust. One could try to adapt the methodology introduced in this work to prove similar tail bounds in other (near-)critical random graphs. We refer the reader to Section 3 for a detailed discussion (and related results) along these lines.
@@
We emphasize that, in Theorem 1.1, $A$ is allowed to depend on $n$. The condition $A=O(n^{1/6})$, which we require in our upper bound for the $\mathbb{G}(n,d,p)$ random graph could be changed to $A=o(n^{2/3})$ at the cost of adding a term of the type $\exp(-\Omega(A^{1/2}n^{1/6}))$ to (6). Moreover, the condition $A=o(n^{2/3})$ which we require in our upper bound for the $\mathbb{G}(n,p)$ model could be relaxed to $A\leq\varepsilon n^{2/3}$ for a sufficiently small constant $\varepsilon>0$. However, when $A=\Theta(n^{2/3})$, then $n^{2/3}/A$ is of constant order and so we would be bounding from above the probability that the size of a maximal cluster is at most some constant, which is not very interesting. Therefore we can say that our upper bound in (5) concerns all the relevant values of $A$. Indeed, we believe the bounds in (5) and (6) to be more interesting for ‘small’ (possibly dependent of $n$) values of $A$, being $n^{2/3}$ the correct order for the size of a largest cluster in both (near-)critical models. We also remark that the parameter $\lambda$, which we consider as fixed, could be allowed to depend on $n$ at the expense of more careful computations. In this case, one would need to keep track of the dependence between $A$ and $\lambda$. Furthermore, if needed, one could give bounds on the ‘smallest values’ of $A_{0}$ for which (5) and (6) hold; similarly, one might try to give bounds on the smallest $n$ for which (5) and (6) are verified (as it is done e.g. in [28]). However, we refrained to do so in order to keep the computations shorter.
@@ 1.2 Related work
They also considered the near-critical model where $p=(1+\lambda n^{-1/3})/n$ for fixed $\lambda\in\mathbb{R}$ and established a similar exponential upper bound.
@@
Amongst other results (some of which were discussed earlier), Nachmias and Peres [27] proved that, in the near-critical regime where $p=(1+\lambda n^{-1/3})/(d-1)$ and $\lambda\in\mathbb{R},3\leq d\in\mathbb{N}$ are both fixed, there are positive constants $c=c(\lambda,d)$ and $C=C(\lambda,d)$ which depend on both $\lambda$ and $d$ such that, for any $A>0$ and for all large enough $n$ (with $dn$ even),
@@
With the purpose of introducing a robust probabilistic methodology capable of recovering the correct asymptotic of Pittel displayed in (7), more recently the author and Roberts [12] introduced a novel method, based on (probabilistic) tools such as generalized ballot theorems [1] and strong embedding with Brownian motion [8], to derive matching upper and lower bounds for the probability of observing unusually large maximal clusters in critical random graphs.
In particular, for the near-critical $\mathbb{G}(n,p)$ model with $p=(1+\lambda n^{-1/3})/n$ and $\lambda\in\mathbb{R}$ (possibly dependent on $n$), the authors showed in [12] that there exist constants $A_{0}>0,n_{0}\in\mathbb{N}$ such that, if $A_{0}\leq A=A(n)=o(n^{1/30}),n\geq n_{0}$ and $\lambda=\lambda(n)$ satisfies $|\lambda|\leq A/3$, then
$\mathbb{P}(|\mathcal{C}_{\max}(\mathbb{G}(n,p))|>An^{2/3})=\Theta\Big{(}\frac{1}{A^{3/2}}\exp\big{(}-\frac{A^{3}}{8}+\frac{\lambda A^{2}}{2}-\frac{\lambda^{2}A}{2}\big{)}\Big{)},$ (10)
thus improving upon the Nachmias and Peres bound given in (8) and retrieving the asymptotic result of Pittel displayed in (7).
Subsequently, the methodology which lead to (10) was adapted by the same authors to study the near-critical $\mathbb{G}(n,d,p)$ random graph where $p=(1+\lambda n^{-1/3})/(d-1)$ with $\lambda\in\mathbb{R}$ (possibly dependent on $n$) and $d\geq 3$ fixed. For this model, it was shown in [13] that there exist constants $A_{0}>0,n_{0}\in\mathbb{N}$ such that, if $A_{0}\leq A=A(n)=o(n^{1/30}),n\geq n_{0}$ and $\lambda=\lambda(n)$ satisfies $|\lambda|\leq A(1-2/d)[3(d-1)]^{-1}$, then
$\mathbb{P}(|\mathcal{C}_{\max}(\mathbb{G}(n,d,p))|>An^{2/3})=\Theta\Big{(}\frac{1}{A^{3/2}}\exp\big{(}-\frac{A^{3}(d-1)(d-2)}{8d^{2}}+\frac{\lambda A^{2}(d-1)}{2d}-\frac{\lambda^{2}A(d-1)}{2(d-2)}\big{)}\Big{)},$
@@
Observe that, in the particular case where $\lambda=0$ (that is, in the purely critical regime), the constant $c(d)\coloneqq(d-1)(d-2)/d^{2}$ which multiplies the leading term $A^{3}/8$ in the exponential above tends to $1$ as $d\rightarrow\infty$. We suspect that this phenomenon is true more generally. Specifically, let $\mathbb{G}$ be a random graph that, when considered at criticality of its parameter(s), exhibits maximal components containing $\Theta(n^{2/3})$ vertices. We believe that, at criticality, there exist constants $c_{1},c_{2}>0$ with $c_{1}<c_{2}$ such that, for all large enough $n$ and (not ‘too large’) $A=A(n)$, then
$\mathbb{P}(|\mathcal{C}_{\max}(\mathbb{G})|>An^{2/3})=\Theta\big{(}A^{-3/2}e^{-C^{*}\frac{A^{3}}{8}}\big{)},$
@@
1.3 Notation and structure of the paper
Notation.
We let =$\mathbb{N}=\{1,2,\dots,\}$ and define ${\mathbb{N}}_{0}\coloneqq{\mathbb{N}}\cup\{0\}$. For $n\in\mathbb{N}$ we set $[n]\coloneqq\{1,2,\ldots,n\}$. Given two sequences of positive real numbers $(x_{n})_{n\geq 1}$ and $(y_{n})_{n\geq 1}$ we write: $(1)$ $x_{n}=O(y_{n})$ if there exist $N\in\mathbb{N}$ and $C\in(0,\infty)$ such that $x_{n}\leq Cy_{n}$ for all $n\geq N$; (2) either $x_{n}=o(y_{n})$ or $x_{n}\ll y_{n}$ if $x_{n}/y_{n}\rightarrow 0$ as $n\rightarrow\infty$; (3) $x_{n}=\Omega(y_{n})$ if $y_{n}=O(x_{n})$; and (4) $x_{n}=\Theta(y_{n})$ if $x_{n}=O(y_{n})$ and $x_{n}=\Omega(y_{n})$ (in which case $y_{n}=\Theta(x_{n})$ too). We often write $c,C$ to denote constants in $(0,\infty)$, and use these letters many times in a single proof even though their actual values may change from line to line.
When talking about random variables, the abbreviation iid means independent and identically distributed. Moreover, given random variables $X$ and $Y$, we write $X\leq_{sd}Y$ if $\mathbb{P}(X\geq z)\leq\mathbb{P}(Y\geq z)$ for every $z\in\mathbb{R}$, and say that $X$ is stochastically dominated by $Y$. We write $\text{Bin}(N,q)$ to denote the binomial distribution of parameters $N\in\mathbb{N}_{0},q\in[0,1]$ and $\text{Unif}([0,1])$ to denote the uniform distribution on $[0,1]$. We use the symbol $=_{d}$ to mean equality in distribution.
Let $\mathbb{G}=(V,E)$ be any (undirected, possibly random) multigraph. Given two vertices $u,v\in V$, we write $u\sim v$ if $\{u,v\}\in E$ and say that vertices $u$ and $v$ are neighbors. We often write $uv$ as shorthand for the edge $\{u,v\}$. We write $u\leftrightarrow v$ if there exists a path of occupied edges connecting vertices $u$ and $v$, where we adopt the convention that $v\leftrightarrow v$ for every $v\in V$. We denote by $\mathcal{C}(v)\coloneqq\{u\in V:u\leftrightarrow v\}$ the component containing vertex $v\in V$. We define a largest component $\mathcal{C}_{\max}(\mathbb{G})$ of $\mathbb{G}$ to be some cluster $\mathcal{C}(v)$ for which $|\mathcal{C}(v)|$ is maximal, so that $|\mathcal{C}_{\max}(\mathbb{G})|=\max_{v\in V}|\mathcal{C}(v)|$.
Structure of the paper.
The rest of the paper is organized as follows. In Section 2 we prove Theorem 1.1, whereas in Section 3 we describe how to modify the martingale argument of [28, 27] with the purpose of obtaining stretched exponential estimates in other two critical models.